### v2

Abstract
Classical evolutionary approaches for multiobjective optimization are quite accurate but incur a lot of queries to the objectives; this can be prohibitive when objectives are expensive oracles. A sample-efficient approach to solving multiobjective optimization is via Gaussian process (GP) surrogates and Bayesian optimization (BO). Multiobjective Bayesian optimization (MOBO) involves the construction of an acquisition function which is optimized to acquire new observation candidates sequentially. This “inner” optimization can be hard due to various reasons: acquisition functions being nonconvex, nondifferentiable and/or unavailable in analytical form; batch sampling usually exacerbates these problems and the success of MOBO heavily relies on this inner optimization. This, ultimately, affects their sample efficiency. To overcome these challenges, we propose a Thompson sampling (TS) based approach ($q\texttt{POTS}$). Whereas TS chooses candidates according to the probability that they are optimal, $q\texttt{POTS}$ chooses candidates according to the probability that they are Pareto optimal. Instead of a hard acquisition function optimization, $q\texttt{POTS}~{}$ solves a cheap multiobjective optimization on the GP posteriors with evolutionary approaches. This way we get the best of both worlds: accuracy of evolutionary approaches and sample-efficiency of MOBO. New candidates are chosen on the posterior GP Pareto frontier according to a maximin distance criterion. $q\texttt{POTS}~{}$ is endowed with theoretical guarantees, a natural exploration-exploitation trade-off and, superior accuracy and sample efficiency than its competitors based on synthetic as well as real-world experiments.
1 Introduction
Mathematically, we consider the simultaneous constrained optimization of $K$ objectives
$\begin{split}\max_{\mathbf{x}\in\boldsymbol{\mathcal{X}}}~{}&\{f_{1}(\mathbf{x}),\ldots,f_{K}(\mathbf{x})\}\\ \textnormal{s.t.}~{}&c_{i}(\mathbf{x})=0,~{}i\in\mathcal{E}\\ ~{}&c_{i}(\mathbf{x})\geq 0,~{}i\in\mathcal{I}\end{split}$ (1)
where $\mathbf{x}$ is the design variable, $\boldsymbol{\mathcal{X}}$ is the feasible set, $\mathbf{x}\in\boldsymbol{\mathcal{X}}\subset\mathbb{R}^{d}$, $f_{k}:\boldsymbol{\mathcal{X}}\rightarrow\mathbb{R},~{}\forall k=1,\ldots,K$, and all $f_{k}$’s are expensive zeroth-order oracles (i.e., no derivative information). $c_{i}$ are nonlinear constraints, also expensive zeroth-order oracles, where $\mathcal{E}$ and $\mathcal{I}$ are the sets of indices of equality and inequality constraints, respectively. We are interested in identifying a Pareto optimal set of solutions, which “Pareto dominates” all other points. Let $\boldsymbol{f(\mathbf{x})}=[f_{1}(\mathbf{x}),\ldots,f_{K}(\mathbf{x})]^{\top}$; when a solution $\boldsymbol{f(\mathbf{x})}$ Pareto dominates another solution $\boldsymbol{f(\mathbf{x}^{\prime})}$, then $f_{k}(\mathbf{x})\geq f_{k}(\mathbf{x}^{\prime}),~{}\forall k=1,\ldots,K$ and there exists $k\in[K]$ such that $f_{k}(\mathbf{x})>f_{k}(\mathbf{x}^{\prime})$. We write Pareto dominance as $\boldsymbol{f(\mathbf{x})}\succ\boldsymbol{f(\mathbf{x}^{\prime})}$. The set $\mathcal{Y}^{*}=\{\boldsymbol{f(\mathbf{x})}~{}:~{}\nexists\mathbf{x}^{\prime}\in\boldsymbol{\mathcal{X}}:\boldsymbol{f(\mathbf{x}^{\prime})\succ\boldsymbol{f(\mathbf{x})}}\}$ is called the Pareto frontier and the set $\boldsymbol{\mathcal{X}}^{*}=\{\mathbf{x}\in\boldsymbol{\mathcal{X}}~{}:~{}\boldsymbol{f(\mathbf{x})}\in\mathcal{Y}^{*}\}$ is called the Pareto set.
We are interested in sample-efficient (that is, minimal queries to the oracles) approaches to solving our problem. Evolutionary algorithms such as the nondominated sorting genetic algorithm (NSGA) Deb et al. (2002) are a popular choice for solving multiobjective problems (see Zitzler et al. (2000) for a review). However, EA based approaches are known to be not sample efficient, which could be prohibitive in real-world applications with expensive oracles. A common approach to solving this problem under such circumstances is to use Bayesian optimization (BO) with Gaussian process (GP) surrogate models for each objective. The idea is to fit a GP model for the objectives using some initial observations $\boldsymbol{\mathcal{D}}^{i}_{n}=\{(\mathbf{x}_{j},y_{j}),~{}j=1,\ldots,n\}$ (where we denote $y_{j}=f_{i}(\mathbf{x}_{j})$) and construct an appropriate acquisition function that quantifies the utility of a candidate point $\mathbf{x}$. Then, an inner optimization problem is solved to choose new point(s) that maximize the acquisition function. The surrogate models are updated with observations at the new points and this process is repeated until a suitable convergence criterion is met. We provide a formal introduction in Section 2.1. BO is very popular in the single objective ($K=1$) case; for instance, the expected improvement (EI) Jones et al. (1998) and probability of improvement (PI) Mockus et al. (1978); Jones (2001), quantifies the probabilistic improvement offered by a candidate point over the “best” observed point so far. The GP upper confidence bound (UCB) Srinivas et al. (2009) makes a conservative estimate of the maximum using the current GP estimate of the function. Other acquisition functions, for $K=1$, include entropy-based Wang and Jegelka (2017); Hernández-Lobato et al. (2014), knowledge gradient (KG) Frazier et al. (2008); Wu and Frazier (2016, 2019), and stepwise uncertainty reduction (SUR) Picheny (2014); Chevalier et al. (2014). Multiobjective Bayesian optimization (MOBO) extends BO to the multiobjective setting; we review some past work in what follows.
1.1 Related work
Naturally, one is tempted to extend the ideas from single objective BO to the multiobjective setting. One of the most common approaches in MOBO are the ones based on scalarizations; that is the $K$ objectives are scalarized into one enabling the use of single objective optimization methods. For instance, the multiobjective efficient global optimization (ParEGO) Knowles (2006) scalarizes the $K$ objectives using randomly drawn weights, which is then treated like a single-objective using EI. Multiobjective evolutionary algorithm based on decomposition (MOEA/D-EGO) Zhang et al. (2009) extends ParEGO to the batch setting using multiple random scalarizations and a genetic algorithm Zhou et al. (2012) to optimize these scalarizations in parallel. Other extensions of ParEGO to the batch setting (qParEGO) and its noisy variant, qNParEGO, have also been proposed Daulton et al. (2021). Paria et al. (2020) do a similar “Chebyshev” scalarization on the objectives but use Thompson sampling (TS) to optimize the scalarized objective. Zhang and Golovin (2020) propose to use a hypervolume scalarization with the property that the expected value of the scalarization over a specific distribution of weights is equivalent to the hypervolume indicator. The authors propose a upper confidence bound algorithm using randomly sampled weights, but provide a very limited empirical evaluation. Scalarizations have the drawback that they are not sample efficient and often fail to capture disconneced and non-convex/concave Pareto frontiers.
Outside of scalarizations, the most common approach in MOBO uses the hypervolume indicator to construct an acquisition function that quantifies the improvement in hypervolume. This usually results in a nonconvex optimization problem that is not straightforward to solve. The EI acquisition function can be extended to the multiobjective setting via the expected hypervolume improvement (EHVI) Couckuyt et al. (2014); Emmerich (2008); gradient based extensions of the EHVI also exist Yang et al. (2019b); Daulton et al. (2020a). Similarly, the SUR criterion extended to multiobjective problems Picheny (2015) is another example. Recent efforts Daulton et al. (2020b, 2021) extend the EHVI idea to sample a batch of acquisitions $(q>1)$ for parallel evaluation and with gradient computation for efficient optimization. However, the stochasticity of the acquisition functions and the associated difficulty in their optimization, particularly as $q$ increases, remains.
Recognizing the local optimal trap of HVI based acquisition functions, several methods emerged which make adhoc changes that better balance exploration and exploitation. Diversity-Guided Efficient Multiobjective Optimization (DGEMO) Konakovic Lukovic et al. (2020) is a recent method for batch MOBO that improves diversity of the chosen points by solving a local optimization problem to maximize hypervolume improvement. Although DGEMO scales well to large batch sizes, it does not account for noisy observations and its ability to account for constraints is unknown. Pareto active learning (PAL) Zuluaga et al. (2013) seeks to classify each candidate point in terms of their probability of being Pareto optimal or not, depending upon the GP posterior uncertainty. This approach presents another difficult acquisition function optimization that is handled via a discretized $\boldsymbol{\mathcal{X}}$ in Zuluaga et al. (2013).
Instead of computing hypervolume improvement, the entropy-based approaches compute the uncertainty reduction in the hypervolume due to new candidates. At the time of writing, these acquisition are considered state of the art, and primarily include predictive entropy search (PESMO) Hernández-Lobato et al. and max-value entropy search (MESMO) Belakaria et al. (2019). MESMO improves upon PESMO in terms of scalability and computational cost however, the acquisition function is still not available in closed form and requires sample approximations. Pareto frontier entropy search (PFES) Suzuki et al. (2020) propose a closed-form entropy criterion for MOBO, but they are restricted to stationary GP kernels, via random Fourier features (RFFs). Additionally, joint entropy search multiobjective optimization (JESMO) Hvarfner et al. (2022); Tu et al. (2022) which considers the joint information gain for the optimal set of inputs and outputs, still depends on approximations to the acquisition function.
The use of Thompson sampling (TS) in MOBO is relatively underexplored to the best of our knowledge. Existing methods still depend on the classical approaches—scalarization and hypervolume improvement—and, thus, inherit their limitations. For instance, the Thompson sampling efficient multiobjective optimization (TSEMO) Bradford et al. (2018) considers the Pareto set identified from posterior GP samples, which are then evaluated for maximum hypervolume improvement. The Thompson sampling with Tcheybchev scalarizations (TS-TCH), likewise, depends on scalarizing objective functions.
Existing limitations and our contributions.
In summary, all of the existing approaches suffer from one or more of the following limitations. (i) Leading to a difficult inner optimization problem that can rarely be solved efficiently and accurately with off-the-shelf optimizers, which negatively impacts their sample efficiency (ii) inability to handle constraints, (iii) inability or difficulty handling batch acquisitions, and (iv) inability to handle noisy objectives .
Our contributions are focused on the following key parameters (i) sample efficiency in both the sequential and batch setting, (ii) ability to handle constraints, (iii) ability to handle noisy objectives, and (iv) ease of implementation. We show that our proposed method addresses all of these parameters while convincingly outperforming the state of the art. Our central idea is inspired from TS in single-objective optimization; whereas TS selects candidate points according to the probability that they are optimal, our approach selects points according to the probability that they are Pareto optimal. Therefore, we call our method Pareto optimal Thompson sampling or $q\texttt{POTS}$, where the $q$ stands for "batch" acquisition. $q\texttt{POTS}$ involves solving a cheap multiobjective optimization problem on the GP posterior sample paths. Our acquisitions are sampled from the computed Pareto sets of the continuous GP posterior sample paths such that they maximize the minimum distance to previously observed points—thereby enabling diversity of candidates and well balancing exploration with exploitation. A very useful consequence of our approach is that selecting a batch of $q$ acquisitions costs the same as the sequential acquisitions. Overall, simplicity of our methodology is our biggest strength.
(a) $n=20$
(b) $n=50$
(c) $n=900$
(d) $n=20$
(e) $n=50$
(f) $n=900$
Figure 1: Demonstration on the Branin-Currin ($d=2,K=2$) test function. Maroon: training points, blue circles: GP posterior samples, blue square: $q\texttt{POTS}$ acquisition, black: true Pareto frontier/set. The input space plots show how $q\texttt{POTS}$ explores and exploits the space (based on distance to maroon). The output space plots show how $q\texttt{POTS}$ asymptotically converges to the solution.
The rationale behind our approach is as follows. Under low posterior uncertainty, the random Pareto frontier is likely a good approximation for $\mathcal{Y}^{*}$, and thus we can exploit. On the other hand, under high posterior uncertainty, the random Pareto frontier is likely far from $\mathcal{Y}^{*}$ and thus we can explore. Therefore our approach is endowed with a natural balance between exploration and exploitation; see Figure 1. The work in the literature that is somewhat similar to us is that of Bradford et al. (2018) (TSEMO) but is significantly different. First, our method circumvents the hypervolume computation part, which scales exponentially with the number of objectives. Second, TSEMO does not support handling constraints, but that is straightforward in $q\texttt{POTS}$. Third, the ability for TSEMO to handle noisy objectives is unknown, but $q\texttt{POTS}$ handles it in a straightforward manner. Fourth, batch acquisition in TSEMO becomes intractable with batch size; in contrast, batch acquisition in our approach comes at the same cost as sequential sampling. Fifth, TSEMO is restricted to stationary GP kernels due to their dependence on random Fourier features Rahimi and Recht (2007); however our approach applies more widely and seamlessly works with nonstationary GP kernels such as deep GPs Damianou and Lawrence (2013); Booth et al. (2023); Sauer et al. (2023); Rajaram et al. (2020) as well as multitask GPs Bonilla et al. (2007); Swersky et al. (2013). Finally, we show how our approach is better scalable in high dimensions by leveraging a Nyström approximation on the posterior covariance matrix. We show several experiments where $q\texttt{POTS}$ convincingly outperforms TSEMO, in addition to other methods in the state of the art. Figure 1 provides a demonstration of $q\texttt{POTS}$ on the Branin-Currin test problem.
The rest of the article is organized as follows. Details of the method are outlined in Section 2, theoretical properties are discussed in Section 3. The numerical experiments are presented in Section 4. We provide concluding remarks and an outlook for future work in Section 5.
2 Methodology
2.1 Single objective Bayesian optimization
We place a GP prior on the oracle $f(\mathbf{x})\sim\mathcal{GP}(0,k(\mathbf{x},\cdot))$, where $k(\cdot,\cdot):\boldsymbol{\mathcal{X}}\times\boldsymbol{\mathcal{X}}\rightarrow\mathbb{R}_{+}$ is a covariance function (a.k.a., kernel). We denote the observations from the oracle as $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i},~{}i=1,\ldots,n$, where we assume that $\epsilon_{i}$ is Gaussian with zero-mean and unknown variance $\tau^{2}$. We begin by fitting a posterior GP (parametrized by hyperparameters $\boldsymbol{\Omega}$) for the observations $\boldsymbol{\mathcal{D}}_{n}=\{\mathbf{x}_{i},y_{i},\tau^{2}\},~{}i=1,\ldots,n$, which gives the conditional aposteriori distribution Rasmussen and Williams (2006).
$\begin{split}Y(\mathbf{x})|\boldsymbol{\mathcal{D}}_{n},\boldsymbol{\Omega}&\sim\mathcal{GP}(\mu_{n}(\mathbf{x}),{\sigma}^{2}_{n}(\mathbf{x})),\\ \mu_{n}(\mathbf{x})&=\mathbf{k}_{n}^{\top}[\mathbf{K}_{n}+\tau^{2}\mathbf{I}]^{-1}\mathbf{y}_{n}\\ \sigma^{2}_{n}(\mathbf{x})&=k(\mathbf{x},\mathbf{x})-\mathbf{k}_{n}^{\top}[\mathbf{K}_{n}+\tau^{2}\mathbf{I}]^{-1}\mathbf{k}_{n},\end{split}$ (2)
where $\mathbf{k}_{n}$ is a vector of covariances between $\mathbf{x}$ and all observed points in $\boldsymbol{\mathcal{D}}_{n}$, $\mathbf{K}_{n}$ is a sample covariance matrix of observed points in $\boldsymbol{\mathcal{D}}_{n}$, $\mathbf{I}$ is the identity matrix, and $\mathbf{y}_{n}$ is the vector of all observations in $\boldsymbol{\mathcal{D}}_{n}$. We also denote by $X_{n}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times d}$ the observation sites. BO seeks to make sequential decisions based on a probabilistic utility function constructed out of the posterior GP $u(\mathbf{x})=u(Y(\mathbf{x})|\boldsymbol{\mathcal{D}}_{n})$. Sequential decisions are the result of an “inner” optimization subproblem that optimizes an acquisition function, typically, of the form $\alpha(\mathbf{x})=\mathbb{E}_{Y|\boldsymbol{\mathcal{D}}_{n}}[u(Y(\mathbf{x}))]$.
An approach that circumvents the acquisition function construction is Thompson sampling (TS) Thompson (1933). TS proposes making a decision according to the probability that it is optimal in some sense Russo and Van Roy (2014). In (single objective) optimization, points are chosen according to the distribution $p_{\mathbf{x}^{*}}(\mathbf{x})$, where $\mathbf{x}^{*}$ is the optimizer of the objective function. In the context of BO, this can be expressed as
[…]
where $\delta$ is the Dirac delta function. Sampling from $p_{\mathbf{x}^{*}}(\mathbf{x})$ is equivalent to finding the maximizer of a sample path $Y(\cdot,\omega)$, where $\omega$ is the random parameter; therefore, in TS for BO, the acquisition function is simply the posterior sample path $\alpha\equiv Y$. Our proposed approach builds on TS which we present next.
2.2 Multiobjective Bayesian optimization
We first fit $K$ independent posterior GP models: $\mathcal{M}_{1},\ldots,\mathcal{M}_{K}$ for the $K$ objectives, with observations $\boldsymbol{\mathcal{D}}^{1:K}_{n}\triangleq\{\boldsymbol{\mathcal{D}}_{n}^{1},\ldots,\boldsymbol{\mathcal{D}}_{n}^{K}\}$. In the multiobjective setting, acquisition functions are typically defined in terms of a hypervolume (HV) indicator. HV of a Pareto frontier $\mathcal{Y}^{*}$ is the $K-$dimensional Lebesgue measure $\lambda$ of the region dominated by $\mathcal{Y}^{*}$ and bounded from below by a reference point $\mathbf{r}\in\mathbb{R}^{K}$. That is, $HV(\mathcal{Y}^{*}|\mathbf{r})=\lambda_{K}(\bigcup_{\nu\in\mathcal{Y}^{*}}[\mathbf{r},\nu])$, where $[\mathbf{r},\nu]$ denotes the hyper-rectangle bounded by vertices $\mathbf{r}$ and $\nu$; see Daulton et al. (2021). EHVI Couckuyt et al. (2014); Emmerich (2008) computes the expectation (with respect to the posterior GP) of the improvement in HV due to a candidate point. Our approach, however, does not involve the computation of HV but is an extension of Equation 3 to the multiobjective setting.
2.3 $q$POTS: Batch Pareto optimal Thompson sampling for multiobjective BO
Inspired from TS, $q\texttt{POTS}$ chooses points according to the probability that they are Pareto optimal. That is, we choose $\mathbf{x}$ according to the probability
$\begin{split}p_{\boldsymbol{\mathcal{X}}^{*}}(\mathbf{x})=\int&\delta\left(\mathbf{x}-arg\max_{\mathbf{x}\in\boldsymbol{\mathcal{X}}}~{}\{Y_{1}(\mathbf{x}),\ldots,Y_{K}(\mathbf{x})\}\right)p(Y_{1}|\boldsymbol{\mathcal{D}}^{1}_{n})\ldots p(Y_{K}|\boldsymbol{\mathcal{D}}^{K}_{n})dY_{1}\ldots dY_{K},\end{split}$ (4)
where the integral is over the support of the joint distribution of all posterior GPs. In practice, this is equivalent to drawing sample paths from the $K$ posterior GPs $Y_{i}(\cdot,\omega),~{}i=1,\ldots,K$, and choosing the next point(s) from its Pareto set $\mathbf{x}_{n+1}\in\operatorname*{arg,max}_{\mathbf{x}\in\boldsymbol{\mathcal{X}}}\{Y_{1}(\mathbf{x},\omega),\ldots,Y_{K}(\mathbf{x},\omega)\}$. We use the “reparametrization trick” to sample from the posterior (again we exclude notations for individual objectives): $Y(\cdot,\omega)=\mu_{n}(\cdot)+{\mathbf{\Sigma}_{n}}^{1/2}(\cdot)\times Z(\omega)$, where $Z$ is a standard multivariate normal random variable. Then, we solve the following cheap multiobjective optimization problem
$X^{*}=\operatorname*{arg,max}_{\mathbf{x}\in\boldsymbol{\mathcal{X}}}\{Y_{1}(\mathbf{x},\omega),\ldots,Y_{K}(\mathbf{x},\omega)\},$ (5)
that can be solved using any method; we choose evolutionary approaches, e.g., the NSGA-II Deb et al. (2002), so that we keep the best of both worlds: accuracy of evolutionary approaches and sample efficiency of MOBO. The next batch of $q$ points is then chosen as $\mathbf{x}_{n+1:q}\in X^{*}.$ That is, we select points from the predicted Pareto set of the sample paths of each posterior GP. In principle, any acquisition chosen from $X^{*}$ is Pareto optimal per the drawn sample paths. However, to promote diversity of candidates, we ensure that the acquisition $\mathbf{x}_{n+1}$ has the largest maximin distance Johnson et al. (1990); Sun et al. (2019) to the previously observed points. This way, there is an automatic balance between exploration and exploitation–see Figure 1 for an illustration. Let $\gamma(\cdot,\cdot):\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ denote the Euclidean distance between two points in $\boldsymbol{\mathcal{X}}$. Then, we pick a batch of $q$ new acquisitions according to the following sequential maximin optimization problems
$\begin{split}\mathbf{x}_{n+1}=&\operatorname*{arg,max}_{{\mathbf{x}}^{*}\in X^{*}}\min_{\mathbf{x}_{i}\in X_{n}}\gamma({\mathbf{x}}^{*},\mathbf{x}_{i})\\ \mathbf{x}_{n+2}=&\operatorname*{arg,max}_{{\mathbf{x}}^{*}\in X^{*}}\min_{\mathbf{x}_{i}\in X_{n}\bigcup\mathbf{x}_{n+1}}\gamma({\mathbf{x}}^{*},\mathbf{x}_{i})\\ \vdots&\\ \mathbf{x}_{n+q}=&\operatorname*{arg,max}_{{\mathbf{x}}^{*}\in X^{*}}\min_{\mathbf{x}_{i}\in X_{n}\bigcup\{\mathbf{x}_{n+1},\ldots,\mathbf{x}_{n+q-1}\}}\gamma({\mathbf{x}}^{*},\mathbf{x}_{i}).\end{split}$ (6)
(6) is a seemingly hard optimization problem, but can be solved quite easily and efficiently. This is because, in practice, the solution of (5) returned by an evolutionary algorithm based optimizer is a point set with finite cardinality. Therefore, what we end up with is the computation of an $N^{*}\times n$ pairwise distance matrix ($N^{*}=|X^{*}|$) and rank-ordering them to according to their column-wise minima. An ascending sort of the resulting vector allows for selecting the $q$ batch acquisition as the last $q$ elements of the sorted vector.
2.4 Extension to constrained problems
Our method seamlessly extends to constrained multiobjective problems and the key step is explained as follows. Let there be a total of $C$ constraints $\{c_{1},\ldots,c_{C}\}$, and $\mathcal{I}\cup\mathcal{E}=[C]$. Further, let $Y_{k+i\in\mathcal{I}}$ and $Y_{k+i\in\mathcal{E}}$ denote the posterior GP of the $i$th inequality constraint and equality constraint functions, respectively. This way, we fit a total of $K+C$ posterior GPs $Y_{1},\ldots,Y_{K},Y_{K+1},\ldots,Y_{K+C}$. For all $i\in[C]$, we define the constrained $q\texttt{POTS}$  to choose points according to
$\begin{split}p_{\boldsymbol{\mathcal{X}}^{*}}(\mathbf{x})=\int\delta&\left(\mathbf{x}-\operatorname*{arg,max}_{\mathbf{x}\in\boldsymbol{\mathcal{X}}}\left[\{Y_{1}(\mathbf{x}),\ldots,Y_{K}(\mathbf{x})\}\times\mathbbm{1}_{\{\bigcap_{i}Y_{k+i\in\mathcal{E}}(\mathbf{x})=0\}}\times\mathbbm{1}_{\{\bigcap_{i}Y_{k+i\in\mathcal{I}}(\mathbf{x})\geq 0\}}\right]\right)\\ &p(Y_{1}|\boldsymbol{\mathcal{D}}_{n})\ldots p(Y_{K}|\boldsymbol{\mathcal{D}}_{n})dY_{1}\ldots dY_{K},\end{split}$ (7)
where $\mathbbm{1}_{\bigcap_{i}Y_{k+i\in\mathcal{I}}(\mathbf{x})\geq 0}$ and $\mathbbm{1}_{\bigcap_{i}Y_{k+i\in\mathcal{I}}(\mathbf{x})=0}$ are the indicator functions for joint feasibility with inequality and equality constraints, respectively, that take a value of $1$ when feasible and 00 otherwise. This is equivalent to drawing samples from the GP posteriors, filtering them out by the indicator functions, and finding a Pareto optimal solution set per the posterior GP sample paths. Then, $q$ maximin candidates are chosen similar to what is done in the unconstrained case. This way, handling constraints in $q\texttt{POTS}$ fits seamlessly into the unconcstrained framework with minimal changes. An illustration on the constrained Branin-Currin test case is shown in the Appendix.
2.5 Nyström approximation on active candidate set
We illustrate how our approach can improve scalability in high dimensions with the Nyström approximation with GP posteriors. We begin by writing the reparametrization technique for a GP posterior:
$\begin{split}Y(X)=&\mu(X)+\left[k(X,X_{n})k(X_{n},X_{n})^{-1}k(X_{n},X)\right]^{1/2}Z\\ =&\mu(X)+\Sigma(X)^{1/2}Z,\end{split}$
where $\Sigma(X)=k(X,X_{n})k(X_{n},X_{n})^{-1}k(X_{n},X)$ is an $N\times N$ posterior covariance matrix, that we also write as $\Sigma_{NN}$; the main computational bottleneck is in computing the square-root of $\Sigma(X)$. We now choose a subset $X_{m}\subset X$, $m\ll N$, to write the following, augmented $(N+m)\times(N+m)$ matrix
$\Sigma_{(N+m)(N+m)}=\begin{bmatrix}\Sigma_{mm}&\Sigma_{mN}\\ \Sigma_{Nm}&\Sigma_{NN}\end{bmatrix},$
where $\Sigma_{mm}$ is the intersection of the $m$ rows and $m$ columns of $\Sigma_{NN}$, corresponding to $X_{m}$, and $\Sigma_{mN}=\Sigma_{Nm}^{\top}$ are the $m$ rows of $\Sigma_{NN}$ corresponding to $X_{m}$. Then, we have that
$\begin{split}\Sigma(X)=&\Sigma_{NN}\approx\Sigma_{mN}\Sigma_{mm}^{-1}\Sigma_{mN}^{\top}\\ =&\left(\Sigma_{mN}\Sigma_{mm}^{-1/2}\right)\left(\Sigma_{mN}\Sigma_{mm}^{-1/2}\right)^{\top}.\end{split}$ (8)
Therefore, we have the Nyström approximation of the posterior covariance matrix square root as $\Sigma(X)^{1/2}\approx\left(\Sigma_{mN}\Sigma_{mm}^{-1/2}\right).$ The approximation in Equation 8 requires an $m\times m$ matrix square root which costs $\mathcal{O}(m^{3})$ and an additional $\mathcal{O}(Nm^{2})$ for the matrix multiplication. Therefore, an overall cost of $\mathcal{O}(m^{3}+Nm^{2})$ is much better than $\mathcal{O}(N^{3})$, provided $m\ll N$. The choice of $X_{m}$ is an open research question; in this work, we choose them to be subset of nondominated points in $\{X_{n},\mathbf{y}_{n}\}$ which works uniformly well in practice. We provide more illustration in the Appendix.
1  Input: With data sets $\boldsymbol{\mathcal{D}}_{n}^{1:K}$, fit $K$ GP models $\mathcal{M}_{1}$ through $\mathcal{M}_{K}$ and GP hyperparameters $\boldsymbol{\Omega}_{1},\ldots,\boldsymbol{\Omega}_{K}$.
2Parameters $B$ (total budget)
  Output: Pareto optimal solution $\mathcal{X}^{*},~{}\mathcal{Y}^{*}$. and fron
  for $i=n+1,\ldots,B$,  do
     1. Sample the posterior GP sample paths: $Y_{i}(\mathbf{x},\omega),~{}\forall i=1,\ldots,K$.
     2. (Optional) Use Nyström approximation for scalability in higher dimensions.
     3. Solve the cheap multiobjective optimization problem via NSGA-II Equation 5.
     4. Choose $q$ candidate points from $X^{*}$ according to (6).
     5. Observe oracles and constraints and append to data set $\boldsymbol{\mathcal{D}}^{1:K}_{i}$.
     6. Update GP hyperparameters $\boldsymbol{\Omega}_{1},\ldots,\boldsymbol{\Omega}_{K}$
  end for
Algorithm 1 $q\texttt{POTS}$: Batch Pareto optimal Thompson sampling
3 Theoretical properties
We now provide theoretical guarantees on $q\texttt{POTS}$, with the following approach. We show that the absolute difference between $f_{i}$ and its GP posterior mean, $\forall\mathbf{x}\in\boldsymbol{\mathcal{X}}$, is asymptotically consistent. Then, we show that the predicted Pareto frontier asymptotically converges to the true Pareto frontier. Our proofs depend on two key assumptions. First, we place regularity assumptions on $f_{i},~{}\forall i$. Then, we assume that our inner (multiobjective) optimization solution, via evolutionary techniques, is solved exactly. We now proceed to present the key theoretical results.
Assumption 3.1 (Exact knowledge of $X^{*}$).
We assume that the inner multiobjective optimization problem $\operatorname*{arg,max}_{\mathbf{x}\in\boldsymbol{\mathcal{X}}}\{Y_{1}(\mathbf{x},\omega),\ldots,Y_{K}(\mathbf{x},\omega)\},$ is solved exactly and the true $X^{*}$ is known.
Assumption 3.2 (Lipschitz).
We assume that the sample paths drawn from the GP prior on $f$ are twice continuously differentiable. Such a GP is achieved by choosing the covariance kernel to be the squared-exponential or from the Matérn class with $\nu=5/2$  (Ghosal and Roy, 2006, Theorem 5). Then, the following holds (for $L\in\mathbb{R}_{+}$): $\mathbb{P}r\left[|f(\mathbf{x})-f(\mathbf{x}^{\prime})|>L\|\mathbf{x}-\mathbf{x}^{\prime}\|\right]\leq e^{-L^{2}/2}.$
Lemma 3.3.
Let Assumption 3.2 hold. Let $\{\mathbf{x}_{i}\},~{}i=1,\ldots,n$ be a sequence of points selected via the $q\texttt{POTS}$ method according to (6). Then, $\lim_{n\rightarrow\infty}~{}\sup_{\mathbf{x}\in\boldsymbol{\mathcal{X}}}\sigma^{2}_{n}(\mathbf{x})=0$.
Using the aforementioned assumptions and Lemma 3.3, we first show that the vector of posterior GPs $\mathbf{Y}=[Y_{1},\ldots,Y_{K}]^{\top}$ converges to the vector of true objectives $\mathbf{f}=[f_{1},\ldots,f_{K}]^{\top}$ with high probability.
Theorem 3.4 (Convergence with high probability).
Let $\delta\in(0,1)$ and Assumption 3.2 hold with $L^{2}=2\log\left(\frac{2}{\delta}\right)$. Let $\tilde{\boldsymbol{\mathcal{X}}}$ denote a discretization of $\boldsymbol{\mathcal{X}}$ with $|\tilde{\boldsymbol{\mathcal{X}}}|<\infty$, where we evaluate the oracles $f_{i}$. Let $\epsilon\geq 0$, and $a_{n}=\sqrt{2\log(b_{n}|\tilde{\boldsymbol{\mathcal{X}}}|)/\delta}$, where $\sum_{n=1}^{\infty}1/b_{n}=1$. Then the vector of posterior GP objectives converges to $\mathbf{f}$ with high probability; that is, the following holds:
$\mathbb{P}r\left[\forall\mathbf{x}\in\boldsymbol{\mathcal{X}},~{}\lim_{n\rightarrow\infty}~{}\|\mathbf{f}(\mathbf{x})-\mathbf{Y}^{n-1}(\mathbf{x})\|_{\infty}\leq\epsilon\right]>1-\delta.$
Due to Theorem 3.4, the Pareto frontier computed with the GP posteriors converges to true Pareto frontier with high probability. All proofs are presented in the Appendix.
4 Experiments
We demonstrate our methodology on several synthetic and real-world experiments varying in input ($d$) and output ($K$) dimensionality. We compare $q\texttt{POTS}$ against TSEMO (our main competitor), PESMO, JESMO, MESMO (state of the art), qNEHVI, qPAREGO (classical methods), and random (Sobol) sampling. Note that, amongst all competitors, only qNPAREGO and qNEHVI support constraints.
Synthetic experiments. For the synthetic experiments, we entertain the Branin-Currin, ZDT3 which has disconnected Pareto frontiers, DTLZ3, and DTLZ7 test functions. Additionally, we test on the constrained OSY problem Osyczka and Kundu (1995) with $C=6$ constraints. Further experiments are included in the Appendix.
Real-world experiments. For the real-world experiments, we use the Vehicle Safety problem Tanabe and Ishibuchi (2020) and the Penicillin production problem Liang and Lai (2021), and the Car side-impact problem Jain and Deb (2013). The Penicillin production problem ($d=7,K=3$) seeks to maximize Penicillin yield while minimizing the time for production as well as the carbon dioxide byproduct emission. The $d=7$ input variables include the culture medium volume, biomass concentration, temperature (in K), glucose substrate concentration, substrate feed rate, substrate feed concentration and the $H+$ concentration Liang and Lai (2021). The Vehicle Safety problem ($d=5,K=3$Tanabe and Ishibuchi (2020) concerns the design of automobiles for enhanced safety. Specifically, we consider the minimization of vehicle mass, the vehicle acceleration during a full-frontal crash, and the toe-board intrusion—a measure of the mechanical damage to the vehicle during an off-frontal crash, with respect to $5$ design variables which represent the reinforced parts of the frontal frame; further detail are found in Liao et al. (2008). The Car side-impact problem is aimed at minimizing the weight of the car while minimizing the pubic force experienced by a passenger and the average velocity of the V-pillar responsible for withstanding the impact load. Additionally, as the fourth objective, the maximum constraint violation of $10$ constraints which include limiting values of abdomen load, pubic force, velocity of V-pillar, and rib deflection is also added Jain and Deb (2013).
We provide each experiment a set of $n=10\times d$ seed samples and observations to start the algorithm, chosen uniformly at random from $\mathcal{X}$; then they are repeated $10$ times. However, the seed is fixed for each repetition across all acquisition policies to keep comparison fair. We add a Gaussian noise with variance $\tau^{2}=10^{-3}$ to all our observations. Our metric for comparison is the hypervolume computed via box decomposition Yang et al. (2019a). Throughout the manuscript, we default to the anisotropic Matern class kernel with $\nu=5/2$. Our implementation primarily builds on GPyTorch Gardner et al. (2018) and BoTorch Balandat et al. (2019); we leverage MPI4Py Dalcin and Fang (2021) to parallelize our repetitions. We use PyMOO Blank and Deb (2020) to leverage its NSGA-II solver; the population size for all problems are fixed as $100\times d$. The full source code to reproduce our results will be provided here333github.com/TBD.
Figure 2 shows the performance comparison for sequential sampling ($q=1$). Notice that $q\texttt{POTS}$ is consistently the best performer, or is amongst the best. We show $q\texttt{POTS}$-Nystrom-Pareto (that is $X_{m}$ chosen as the current Pareto set) for the $d=10$ experiments which shows no loss of accuracy for the DTLZ3 function and, surprisingly, a slight gain in accuracy for the ZDT3 function. In the bottom right, we show the performance on the constrained OSY Osyczka and Kundu (1995) problem; qNEHVI failed to complete for this experiment due to memory overflow. Additional experiments are provided in the Appendix. Despite the success in the $q=1$ setting, our main benefit is seen in the batch sampling setting which we discuss next.
The demonstration for the batch setting is shown in Figure 3; we present $q=2,4$ with more in the Appendix; like in the $q=1$ setting we include $q\texttt{POTS}$-Nystrom-Pareto for the $d=10$ experiments only. While $q\texttt{POTS}$ comfortably wins in all experiments, notice that the difference with TSEMO is more pronounced in the batch setting. We demonstrate the computational efficiency gains of $q\texttt{POTS}$ in the Appendix.
5 Conclusion
We present $q\texttt{POTS}$, a sample efficient method to solve expensive constrained multiobjective optimization problems. Classical evolutionary approaches are accurate but not sample efficient. Bayesian optimization is, in general, sample efficient, but its success relies heavily on solving a difficult inner optimization. Our proposed $q\texttt{POTS}$ relies on sampling from the posterior GPs, and solving a cheap multiobjective optimization on the posterior sample paths with classical evolutionary algorithms such as NSGA-II. This way, we bridge the gap between classical EA’s and MOBO and thus, in a sense, getting the best of both worlds. $q\texttt{POTS}$ naturally balances exploration and exploitation, can handle constraints and noisy objectives, and can efficiently handle batch acquisitions; no other existing method has all these features and still outperform $q\texttt{POTS}$ to the best of our knowledge. Scaling up to higher dimensions is achieved through a Nyström approximation which significantly speeds up the method without any practical loss of accuracy. Crucially, our approach freely applies to any covariance kernel (unlike RFF for stationary kernels).
Looking into the future, one avenue of interest is marrying $q\texttt{POTS}$ with deep GPs which would open doors for more applications where nonstationarity in the oracles is expected to be strong.
Figure 2: Sequential acquisition. Hypervolume Vs. iterations for sequential ($q=1$) acquisition. $q\texttt{POTS}$ outperforms all competitors or is amongst the best. Bottom right shows constrained case on the OSY problem Osyczka and Kundu (1995). Additional experiments are provided in the appendix.
Figure 3: Batch acquisition. Hypervolume Vs. iterations for batch ($q>1$) acquisition. $q\texttt{POTS}$ outperforms all competitors, but the benefit is more pronounced in the batch case. Additional experiments, including constrained problems, are shown in the Appendix.
Appendix
6 ADDITIONAL EXPERIMENTS
(a) qNEHVI
(b) qNPAREGO
(c) Sobol
(d) PESMO
(e) MESMO
(f) $q\texttt{POTS}$ (Ours)
Figure 4: Demonstration on the ZDT3 ($K=2$) test function that has disjoint Pareto frontiers. Navy are the true Pareto frontier and red are the predicted Pareto frontier. With $20$ seed points and only an additional $204$ batch acquisitions at $q=4$, notice that $q\texttt{POTS}$ is quickly able to resolve the true Pareto frontier while other methods in the state of the art struggle to come any close.
[…]
Figure 5: Additional constrained real-world experiments; top row: sequential, bottom row: batch acquisitions. Note that entropy based acquisition functions (PESMO, MESMO, JESMO) and TSEMO do not support constraints. We were not able to get any results with $qNEHVI$ due to memory overflow issues. Therefore, we have compared against qNParEGO and Sobol. Note that we ran qPOTS-Nystrom-Pareto only for $d\geq 10$ cases as it is not necessary for low dimensional cases.
We show the performance on two constrained problems: the OSY problem Osyczka and Kundu (1995) with $2$ objectives and $6$ constraints, and the discbrake Tanabe and Ishibuchi (2020) problem with $2$ objectives and $4$ constraints. Note that, TSEMO as well as the entropy based acquisition functions don’t support constraints. We had trouble running constrained problems with qNEHVI—the code took extraordinarily long time to execute and we had memory overflow issues. We perceive this as a computational bottleneck in qNEHVI. On the other hand, $q\texttt{POTS}$ requires no special modifications for handling constraints and the computational overhead is only due to fitting additional GPs for the constraints. We consider this one of the biggest strengths of $q\texttt{POTS}$.
We also show additional synthetic experiments in Figure 6 which includes high-dimensional experiments up to $d=20$. Again, $q\texttt{POTS}$ is either the best or amongst the best in all cases.
Figure 6: Additional synthetic batch experiments. $q\texttt{POTS}$ is either the best or amongst the best.
Figure 7: Cumulative times for sequential acquisition ($q=1$).
Figure 8: Cumulative times for batch acquisition ($q>1$).
Finally, we show the computational times for all algorithms, including $q\texttt{POTS}$, each of the experiments. Figure 7 and Figure 8 show them for sequential and batch acquisitions, respectively. First, notice that $q\texttt{POTS}$ is much cheaper than the rest, with the exception of Sobol, for low dimensional problems. I higher dimensions, and with more objectives and constraints, $q\texttt{POTS}$ is no worse than the rest. The benefit is even more prominent in the batch acquisition setting where the cost of $q\texttt{POTS}$ remains almost the same as the sequential setting but other acquisition policies become more expensive.
7 MISSING PROOFS
We now provide theoretical guarantees on $q\texttt{POTS}$, with the following approach. We show that the absolute difference between $f_{i}$ and its GP posterior mean, $\forall\mathbf{x}\in\boldsymbol{\mathcal{X}}$, is bounded for finite samples and is asymptotically consistent. Then, we show that the predicted Pareto frontier asymptotically converges to the true Pareto frontier. Our proofs depend on two key assumptions. First, we place regularity assumptions on $f_{i},~{}\forall i$. Then, we assume that our inner (multiobjective) optimization solution, via evolutionary techniques, is solved exactly. We now proceed to present the key theoretical results.
Assumption 7.1 (Exact knowledge of $X^{*}$).
We assume that the inner multiobjective optimization problem $\operatorname*{arg,max}_{\mathbf{x}\in\boldsymbol{\mathcal{X}}}\{Y_{1}(\mathbf{x},\omega),\ldots,Y_{K}(\mathbf{x},\omega)\},$ is solved exactly and the true $X^{*}$ is known. In practice, this assumption is reasonable and, to improve accuracy, can be fine-tuned with larger $\tilde{N}$ and $N_{\textnormal{gen}}$.
Assumption 7.2 (Lipschitz).
We assume that the sample paths drawn from the GP prior on $f$ are twice continuously differentiable. Such a GP is achieved by choosing the covariance kernel to be four times differentiable (Ghosal and Roy, 2006, Theorem 5), for example, the squared-exponential or Matérn-class kernel. Then, the following holds (for $L\in\mathbb{R}_{+}$): $\mathbb{P}r\left[|f(\mathbf{x})-f(\mathbf{x}^{\prime})|>L\|\mathbf{x}-\mathbf{x}^{\prime}\|\right]\leq e^{-L^{2}/2}.$
7.1 Proof of Lemma 3.3
Lemma 7.3.
Let Assumption  7.2 hold. Let $\{\mathbf{x}_{i}\},~{}i=1,\ldots,n$ be a sequence of points selected via the $q\texttt{POTS}$ method according to Eq. (6) with $q=1$. Then, $\lim_{n\rightarrow\infty}~{}\sup_{\mathbf{x}\in\boldsymbol{\mathcal{X}}}\sigma^{2}_{n}(\mathbf{x})=0$.
Proof.
First of all, note that $q\texttt{POTS}$ chooses new acquisitions according to the maximin distance criterion (Eq. (6)). Therefore, unless the posterior uncertainty is 00 almost everywhere, the same acquisition is never chosen more than once.
W.l.o.g, let $k(\mathbf{x},\mathbf{x})=1$. Recall that the posterior variance of the GP is given by
$\sigma_{n}^{2}(\mathbf{x})=k(\mathbf{x},\mathbf{x})-\mathbf{k}_{n}^{\top}\mathbf{K}_{n}^{-1}\mathbf{k}_{n}.$
Note that as $n\rightarrow\infty$, then $\mathbf{k}_{n}\rightarrow\mathbf{K}^{:,i}_{n}$, where by $\mathbf{K}^{:,i}_{n}$ we mean the $i$th column of $\mathbf{K}_{n}$, $i\in[n]$. This is because, when the $f$ has been observed at infinitely many points $\{\mathbf{x}_{i}\}_{i=1}^{\infty}$ and recall that $X_{n}$ is the row stack of all of these points, then $\mathbf{k}_{n}=k(\mathbf{x},X_{n})$ represents the covariance between a candidate point $\mathbf{x}$ and all previous observation sites $X_{n}$. When $n=\infty$, $\mathbf{x}$ is likely very close to one of the points $\mathbf{x}_{i}\in\{\mathbf{x}_{i}\}_{i=1}^{\infty}$ and hence $k(\mathbf{x},X_{n})\rightarrow k(\mathbf{x}_{i},X_{n}),~{}i\in[n]$. And therefore $\mathbf{k}_{n}^{\top}\mathbf{K}_{n}^{-1}=\mathbf{e}_{i}$, where $\mathbf{e}_{i}$ is an $n$-vector with $1$ at the $i$th location and 00 elsewhere. Then
$\lim_{n\rightarrow\infty}\sigma_{n}^{2}(\mathbf{x})=1-\mathbf{K}^{:,i~{}\top}_{n}\mathbf{e}_{i}=1-k(\mathbf{x}_{i},\mathbf{x}_{i})=0.$
Using the aforementioned assumptions and Lemma 7.3, we first show that the vector of posterior GPs $\mathbf{Y}=[Y_{1},\ldots,Y_{K}]^{\top}$ converges to the vector of true objectives $\mathbf{f}=[f_{1},\ldots,f_{K}]^{\top}$ with high probability.
Theorem 7.4 (Convergence with high probability).
Let $\delta\in(0,1)$ and Assumption 3.2 hold with $L^{2}=2\log\left(\frac{2}{\delta}\right)$. Further, let $g$ be some affine function that operates on $f$. Let $\tilde{\boldsymbol{\mathcal{X}}}$ denote a discretization of $\boldsymbol{\mathcal{X}}$ with $|\tilde{\boldsymbol{\mathcal{X}}}|<\infty$, where we evaluate the oracle $f$. Let $\epsilon\geq 0$, and $a_{n}=\sqrt{2\log(b_{n}|\tilde{\boldsymbol{\mathcal{X}}}|)/\delta}$, where $\sum_{n=1}^{\infty}1/b_{n}=1$. Then the vector of posterior GP objectives converges to $\mathbf{f}$ with high probability; that is, the following holds:
$\mathbb{P}r\left[\forall\mathbf{x}\in\boldsymbol{\mathcal{X}},~{}\lim_{n\rightarrow\infty}~{}\|\mathbf{f}(\mathbf{x})-\mathbf{Y}^{n-1}(\mathbf{x})\|_{\infty}\leq\epsilon\right]>1-\delta.$
Proof of Theorem 3.4.
We first present a simple identity that bounds the probability that a Gaussian random variable is greater than some value. Given $r\sim\mathcal{N}(0,1)$ and $c>0$, consider the following:
[…]
where the second line is due to the fact that $0\leq e^{-c(r-c)}\leq 1$ for $r\geq c$ and $c>0$ and the last line from the fact that the integrand is the density of $\mathcal{N}(c,1)$. Using this result and the prior assumption that the unknown function $f$ has a GP prior, we state, with $c=a_{n}$ and a fixed $\mathbf{x}_{i}\in\tilde{X}$,
$\begin{split}\mathbb{P}r\left[|f(\mathbf{x}_{i})-\mu_{n-1}(\mathbf{x}_{i})|>a_{n}\sigma_{n-1}(\mathbf{x}_{i})\right]\leq&\frac{1}{2}e^{-a_{n}^{2}/2}=\frac{\delta}{2b_{n}|\tilde{\mathcal{X}}|}.\end{split}$ (10)
Applying the union bound $\forall i$, we get
$\begin{split}\mathbb{P}r\bigcup_{i=1}^{|\tilde{\mathcal{X}}|}|f(\mathbf{x}_{i})-\mu_{n-1}(\mathbf{x}_{i})|>a_{n}\sigma_{n-1}(\mathbf{x}_{i})&\leq\sum_{i=1}^{|\tilde{\mathcal{X}}|}\mathbb{P}r\left[|f(\mathbf{x}_{i})-\mu_{n-1}(\mathbf{x}_{i})|>a_{n}\sigma_{n-1}(\mathbf{x}_{i})\right]\\ &\leq\sum_{i=1}^{|\tilde{\mathcal{X}}|}\frac{\delta}{2b_{n}|\tilde{\mathcal{X}}|}=\frac{\delta}{2b_{n}}.\end{split}$
Now we apply the union bound $\forall n$:
$\begin{split}\mathbb{P}r\left[\forall\tilde{\mathcal{X}}\in\tilde{\mathcal{X}},~{}\bigcup_{n=1}^{\infty}|f(\tilde{\mathcal{X}})-\mu_{n-1}(\tilde{\mathcal{X}})|>a_{n}\sigma_{n-1}(\tilde{\mathcal{X}})\right]\leq&\sum_{n=1}^{\infty}\frac{\delta}{2b_{n}}=\frac{\delta}{2}.\\ \end{split}$ (11)
We would like to extend the proof to a general $\mathbf{x}\in\boldsymbol{\mathcal{X}}$. For some $\mathbf{x}\in\boldsymbol{\mathcal{X}}$, let $\tilde{\mathcal{X}}$ denote a point $\in\tilde{\mathcal{X}}$ that is nearest (in Euclidean distance) to $\mathbf{x}$. Then, per Assumption 3.2 and by setting $L^{2}=2\log\left(\frac{2}{\delta}\right)$, the following statement holds:
$\mathbb{P}r\left[|f(\mathbf{x})-f(\tilde{\mathbf{x}})|>\sqrt{2\log\left(\frac{2}{\delta}\right)}\|\mathbf{x}-\tilde{\mathbf{x}}\|\right]\leq\frac{\delta}{2}.$ (12)
Combining (11) and (12), the following statement holds:
$\mathbb{P}r\left[\forall\mathbf{x}\in\boldsymbol{\mathcal{X}},\forall n,~{}|f(\mathbf{x})-\mu_{n-1}(\tilde{\mathbf{x}})|\leq a_{n}\sigma_{n-1}(\mathbf{x})+\sqrt{2\log\left(\frac{2}{\delta}\right)}\|\mathbf{x}-\tilde{\mathbf{x}}\|\right]>1-\delta.$
From Lemma 3.2, the $\lim_{n\rightarrow\infty}\sup_{\mathbf{x}\in\boldsymbol{\mathcal{X}}}\sigma_{n-1}(\mathbf{x})=0$, and thus
$\mathbb{P}r\left[\forall\mathbf{x}\in\boldsymbol{\mathcal{X}},\lim_{n\rightarrow\infty}|f(\mathbf{x})-\mu_{n-1}(\tilde{\mathbf{x}})|\leq\epsilon\right]>1-\delta,$
where $\epsilon=\sqrt{2\log\left(\frac{2}{\delta}\right)}\|\mathbf{x}-\tilde{\mathbf{x}}\|$. Circling back to the reparametrization trick, the above statement can be restated as
$\begin{split}\mathbb{P}r\left[\forall\mathbf{x}\in\boldsymbol{\mathcal{X}},\lim_{n\rightarrow\infty}|f(\mathbf{x})-Y^{n-1}(\tilde{\mathbf{x}})+\sigma_{n-1}(\mathbf{x})z|\leq\epsilon\right]>&1-\delta\\ \mathbb{P}r\left[\forall\mathbf{x}\in\boldsymbol{\mathcal{X}},\lim_{n\rightarrow\infty}|f(\mathbf{x})-Y^{n-1}(\tilde{\mathbf{x}})|\leq\epsilon\right]>&1-\delta,\end{split}$
where $Y^{n-1}$ is the posterior GP after the $(n-1)$th round, $z$ is the standard normal random variable, and the second line follows from Lemma 3.3. By extension, the above statement holds for all objectives $f_{i},~{}\forall i\in[K]$. Therefore, the above statement can be generalized as
$\mathbb{P}r\left[\forall\mathbf{x}\in\boldsymbol{\mathcal{X}},\lim_{n\rightarrow\infty}\|\mathbf{f}(\mathbf{x})-\mathbf{Y}^{n-1}(\tilde{\mathbf{x}})\|_{\infty}\leq\epsilon\right]>1-\delta.$
Due to Theorem 3.4, the Pareto frontier computed with the GP posteriors converges to true Pareto frontier with high probability. Note that Figure 1f in the paper confirms this empirically.